(d) If the ball was replaced by a ball which had significantly more mass, but the force exerted on the ball matched the original graph shown above, how will this affect the launch speed of the ball? The launch speed of the greater-mass ball will be…

a) Greater b) Smaller c) The same

Justify your response by referencing work and/or energy principles.

(e) With the original ball (m = 0.008 kg), the cannon is changed so that it applies a force of 50.0 N for the entire 4 cm distance, rather than the variable force which peaks at 70.0 N shown originally. With the cannon applying a constant force of 50.0 N, the distance d the ball will travel before stopping will be…

a) Greater b) Smaller c) The same

Justify your response by referencing work and/or energy principles.

1
Expert's answer
2021-03-26T11:40:30-0400

d)


"K=0.5mv^2=const"

Answer: b) Smaller 

e)


"W=0.5(70)(0.04)=1.4\ J\\W'=(50)(0.04)=2\ J\\d'=\frac{W'}{F_f}\\d=\frac{W}{F_f}\\d'>d"

Answer: a) Greater